OpenStudy (anonymous):

how do I simplify 108 times the square root of 3?

OpenStudy (anonymous):

wouldn't that just me \[108\sqrt{3}\] ?

OpenStudy (anonymous):

its\[180\times \sqrt{3}\]

geerky42 (geerky42):

is it 108√3 ? in that case, it cannot be simplified more what it already is.